OpenStudy (anonymous):

3. What number would go in the blanks to create a perfect square trinomial? (x2 + 10x + _______) = _______ + 6 A. 20 B. 10 C. 5 D. 25

OpenStudy (anonymous):

D

OpenStudy (anonymous):

25 is a perfect square, so i think its d.

OpenStudy (anonymous):

4. What number would go in the blanks to create a perfect square trinomial? (x2 - 6x + _______) = _______ + 12 A. 9 B. 6 C. 12 D. 3 @Mr.Man420

OpenStudy (anonymous):

A

OpenStudy (anonymous):

5. Fill in the blank with the correct number to convert 0 = x2 - 4x + 8 into transformational form 0 = (x - ___)2 + 4. A. 4 B. 8 C. 16 D. 2 @Mr.Man420

OpenStudy (anonymous):

Ahh sorry, im not quite sure on this one...

OpenStudy (anonymous):

My pleasure helping you tho(: